OpenStudy (destinyyyy):

A=p(1+rt) p=12000 r=9% t=14

OpenStudy (destinyyyy):

I got the answer A=183,120 ... When I checked it on a website it said its 12000... Just want another opinion. Thank you

OpenStudy (anonymous):

A= 12000(1+(.09*14))

OpenStudy (er.mohd.amir):

A=1200(1+9*14/100) solve it according to Formula written 9%=9/100

OpenStudy (destinyyyy):

12.000(1(0.09)(14) 12000(1.09(14) 12000(15.26) 183,120

OpenStudy (destinyyyy):

I dont see where I solved it wrong. But the math website gives a different answer.. Am I right or wrong?